StarT the Next Generation: Integrating Global Caches and Dataflow

Architecture

Computation Structures Group Memo 354 February 25, 1994

Boon Seong Ang Arvind Derek Chiou

In Proceedings of the International Conference on Computer Systems and Education, IISc, Bangalore, India, April, 1994

This report describes research done at the Laboratory for Computer Science of the Massachusetts Institute of Technology. Funding for the Laboratory is provided in part by the Advanced Research Projects Agency of the Department of Defense under the Office of Naval Research contract N00014-92-J-1310.